OpenStudy (anonymous):

how would you solve 2x+6y=-5. you have to find the equation in slope-intercept form?

OpenStudy (sandra):

well, slope intercept form is y=mx + b, where m is the slope

OpenStudy (sandra):

so what you want to do is get y on one side by itself

OpenStudy (sandra):

so 2x + 5 + 6y = 0

OpenStudy (sandra):

2x + 5 = -6y

OpenStudy (sandra):

divide both sides by -6

OpenStudy (sandra):

-2/6x -5/6 = y

OpenStudy (sandra):

or the other way around: y = -1/3x -5/6

OpenStudy (sandra):

so the slope is 1/3

OpenStudy (sandra):

and that is slope intercept form